• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name averett

    Log entry time 16:04:57 on June10,2009

    Entry number 280419

    keyword=day shift summary

    0800 Production running, transverse target
    0855 NMR=58.2%, BHWP is now OUT, New Compton run started, new HAPPEX run started=31506
    0900 Start production runs, 3438, 22309
    1025 Runs 3438, 3439, 22309, 22310 have the BHWP status incorrectly listed as IN and the target polarization incorrectly listed as 56.7%.
    1140 Target NMR=58.3%, Yi will rotate target field to longitudinal, positive, move to N2 reference cell, pressure=18 psig.
    12:10 Move back to pol3he target position. Orientation is longitudinal, positive. NMR measurement=55.8%
    1325 MCC informs us that the cooling water to the beam dump is tripped off. No Beam until fixed. End HAPPEX and Com[pton runs too.
    14:02 Resume production running